OpenStudy (anonymous):

Find the fourth degree polynomial with real coefficients having the zeros: 0, 1, 1+i

OpenStudy (anonymous):

complex zeros always occur in pairs. so zeros are 0,1,1+i,1-i if a,b,c,d are zeros of a polynomial ,then polynomial is f(x)=(x-0)(x-1)(x-(1+i))(x-(1-i)) solve it.

OpenStudy (anonymous):

I know that much. I just need to know the next step after that.

OpenStudy (anonymous):

What happens to the imaginary number? How do I solve with it?

OpenStudy (anonymous):

\[(x-1-i)(x-1+i)=\left\{ \left( x-1 \right) -i\right\}\left\{ \left( x-1 \right)+i \right\}=\left( x-1 \right)^2-i^2=\left( x-1 \right)^2+1\]

OpenStudy (anonymous):

\[=x^2-2x+1+1=x^2-2x+2\]

OpenStudy (anonymous):

Apparently the correct answer on the paper is x^4-3x^3+4x^2-2x

OpenStudy (anonymous):

\[f \left( x \right)=x(x-1)(x^2-2x-2)\] simplify it.

OpenStudy (anonymous):

I'm so confused as to where you got that. I don't understand this at all.

OpenStudy (sleepyjess):

\(\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\\\huge\cal\color{red}W\color{orange}E\color{goldenrod}L\color{yellow}C\color{lightgreen}O\color{darkgreen}M\color{turquoise}E~~\color{royalblue}T\color{purple}O~~\color{#00bfff}{Open}\color{#11c520}{Study}\\\huge\color{red}{\bigstar}\color{orange}{\bigstar}\color{yellow}{\bigstar}\color{lightgreen}{\bigstar}\color{green}{\bigstar}\color{turquoise}{\bigstar}\color{royalblue}{\bigstar}\color{purple}{\bigstar}~\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#00bfff}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\color{#11c520}{\bigstar}\)